### v1

# Comparing Inferential Strategies of Humans and Large Language Models in Deductive Reasoning

Abstract
Deductive reasoning plays a pivotal role in the formulation of sound and cohesive arguments. It allows individuals to draw conclusions that logically follow, given the truth value of the information provided. Recent progress in the domain of large language models (LLMs) has showcased their capability in executing deductive reasoning tasks. Nonetheless, a significant portion of research primarily assesses the accuracy of LLMs in solving such tasks, often overlooking a deeper analysis of their reasoning behavior. In this study, we draw upon principles from cognitive psychology to examine inferential strategies employed by LLMs, through a detailed evaluation of their responses to propositional logic problems. Our findings indicate that LLMs display reasoning patterns akin to those observed in humans, including strategies like supposition following or chain construction. Moreover, our research demonstrates that the architecture and scale of the model significantly affect its preferred method of reasoning, with more advanced models tending to adopt strategies more frequently than less sophisticated ones. Importantly, we assert that a model’s accuracy, that is the correctness of its final conclusion, does not necessarily reflect the validity of its reasoning process. This distinction underscores the necessity for more nuanced evaluation procedures in the field.
0.5pt
Comparing Inferential Strategies of Humans and Large Language Models in Deductive Reasoning
Philipp Mondorf1, 2 and Barbara Plank1, 2 1MaiNLP, Center for Information and Language Processing, LMU Munich, Germany 2Munich Center for Machine Learning (MCML), Munich, Germany p.mondorf@lmu.de b.plank@lmu.de
1 Introduction
Deductive reasoning, that is the process of drawing conclusions that logically follow from the information at hand, is an integral aspect of human cognition and plays a pivotal role in formulating sound and coherent arguments Leighton (2003). Take, for example, the following statements:
If there is a blue marble in the box then there is a green marble in the box.
There is a blue marble in the box.
Even without proper training in logic, most individuals can naturally deduce the valid conclusion:
Therefore, there is a green marble in the box.
This innate capability of drawing conclusions that invariably follow from the truth value of available information has been a focal point of scholarly interest for centuries Holyoak and Morrison (2005). Propositional logic, a subfield of deductive reasoning, focuses on constructing logical arguments based on the relationship between statements similar to those in the example previously mentioned Hurley (2011). Extensive research has been dedicated to examining human reasoning behavior in contexts that involve propositional logic. For instance, Van der Henst et al. (2002) have identified five different strategies people commonly employ when navigating problems of propositional logic (see Section 2). Such behavioral studies have been crucial in shaping theories that shed light on the fundamental elements of cognitive reasoning processes Rips (1994); Johnson-Laird (1986); Kahneman et al. (1982).
In parallel, recent advancements in the field of large language models have demonstrated their potential in executing tasks involving deductive reasoning Yang et al. (2023); Yu et al. (2023); Huang and Chang (2023). Yet, the extent to which LLMs truly possess such abilities remains a subject of ongoing debate Mahowald et al. (2023); Mitchell and Krakauer (2023). Unlike behavioral studies in human reasoning that are often characterized by in-depth examinations of the reasoners’ expressions, many studies on LLM-based reasoning tend to focus on task performance and accuracy metrics, offering limited insights into the underlying reasoning behavior of the models Mitra et al. (2023); OpenAI et al. (2023); Team et al. (2023).
In this paper, we draw from the cognitive science literature Van der Henst et al. (2002) and study inferential strategies employed by LLMs when solving propositional logic problems (see Figure 1). We analyze the reasoning behavior of three different language model families, varying in model size and fine-tuning procedure, and compare them to the behavior found in humans. To the best of our knowledge, we are the first to comprehensively compare inferential strategies employed by large language models and humans. We analyze the models’ output both quantitatively and qualitatively via manual inspection, to provide insights into the soundness of their verbalized reasoning strategies. Our findings reveal that:
All models exhibit inferential strategies akin to those observed in human reasoning, such as supposition following and chain construction.
The inferential strategy employed is significantly influenced by the model family, as different families favor different approaches.
Models are often right but for the wrong reasons: the accuracy of a model, that is the number of correct final answers, does not reflect whether its reasoning is sound, i.e. logically follows from the statements at hand.
The strategy employed by a model is closely related to the soundness of its reasoning, where certain strategies lead to correct reasoning and others tend to introduce errors.
In contrast to human reasoners, models occasionally adopt a symbolic strategy, where formal logical calculus is employed to solve the problems of propositional logic.
Through this work, we hope to advance the understanding of reasoning in LLMs.
2 Strategies in Propositional Reasoning
Propositional logic studies the relationships among statements (or propositions) and the methods for constructing logical arguments based on them Hurley (2011). At the core of propositional logic are simple statements that can be combined through the use of logical connectives such as "not", "and", "or", and "if… then…", thereby forming more complex compound statements. Conclusions are logically deduced, where the truth value of the propositions necessitates the truth of the conclusion. This form of logical reasoning allows us to construct sound arguments that are invariably true, given the truth value of the information provided. As such, propositional logic is fundamental to various disciplines, including science, mathematics, and philosophy, where it offers a structured approach to reasoning and argumentation.
To gain insights into the inferential processes humans employ in propositional reasoning, Van der Henst et al. (2002) conducted a series of experiments that study the behavior of participants during propositional reasoning. They formulated straightforward propositional logic problems with neutral content (the presence or absence of colored marbles in a box, similar to the problem illustrated in Figure 1) and requested participants to articulate their thought processes while engaging with these problems. Participants were permitted the use of paper and pencil for their workings. Both their verbal explanations and written responses were meticulously recorded, transcribed and analyzed thereafter. Van der Henst et al. (2002) discovered five strategies reasoners commonly utilize to navigate the problems, offering insights into their inferential mechanisms employed in propositional reasoning. In the following, we give a short description of each strategy (illustrated in Figure 2). For more details and additional examples, we refer to the original study by Van der Henst et al. (2002).
Incremental Diagram. This strategy involves the creation of a comprehensive diagram that keeps track of all potential outcomes compatible with the statements of the problem. During the reasoning process, individuals progressively increment their diagrams to incorporate new information derived (see left box in Figure 2). The result is a single diagram that records a variety of possibilities compatible with the premises, often including even those that might be irrelevant to the task.111In contrast to Van der Henst et al. (2002), we in fact observe no single occurrence of the incremental diagram strategy in LLMs, despite the authors finding this strategy being employed most frequently by humans. We belief that this discrepancy stems from the use of pen and paper in human assessments, implicitly encouraging diagrammatic reasoning.
Supposition Following. Reasoners employing this strategy start with a supposition, e.g. by assuming a marble of a certain color. They then trace the consequences of that supposition following from the statements at hand, as illustrated in the second box from the left of Figure 2. The result is a sequence of literals (in this case, marbles of a certain color) without logical connectives. The success of supposition following strongly depends on the supposition made by the reasoner. Some suppositions might lead to inferences that are relevant to the problem, while others might lead to irrelevant conclusions.
Chain Construction. When employing this strategy, reasoners construct a chain of conditional statements derived from the premises at hand or from intermediate deductions. An example of chain construction is displayed in the central box of Figure 2. Premises are converted into a chain of conditional statements that are linked by their entities. A distinctive feature of this chain is the interconnection between conditionals, where the consequent of one conditional is the antecedent of the following.
Compound Strategy. Reasoners following the compound strategy combine two or more statements to derive a new compound conclusion. This process yields a series of novel conclusions, each building upon the preceding ones. An example illustrating this strategy is given in the upper second box from the right of Figure 2. Based on the first two premises, the compound conclusion: “If red then not blue.” is inferred, and then used to draw another compound conclusion (“If red or pink then not blue.”) together with the last premise of the problem statement.
Concatenation Strategy. This approach entails the concatenation of two or more statements into a single conclusion encompassing the logical implications of each combined proposition. This strategy is subtle and has only been infrequently observed by Van der Henst et al. (2002). An example of the strategy is illustrated in the lower second box from the right of Figure 2.
Symbolic Strategy. We could identify an additional strategy occasionally employed by LLMs, which has not been observed by Van der Henst et al. (2002) in human reasoners. This strategy, which we denote as symbolic strategy, is characterized by models employing formal logical calculus to solve the tasks at hand. When following this strategy, models either translate logical statements that are expressed in natural language (e.g. “If there is a white marble then there is not a red marble.”) into formal logic (), and then operate on those expressions, or create a truth table from which they aim to infer the validity of the conclusion. An illustration of this strategy is provided in the right box of Figure 2.
3 Experimental Setup
Task Overview. Our task setup aligns with the experiment conducted by Van der Henst et al. (2002) to allow for a fair comparison between the inferential strategies found in humans and those identified in LLMs.222More specifically, experiment one of Van der Henst et al. (2002). In particular, we evaluate each model on the 12 problems of propositional logic suggested by Van der Henst et al. (2002) (an overview of each problem can be found in Figure 5 in the appendix). In each problem, models are presented with a set of statements (or premises) and must determine whether a given conclusion logically follows (for an example, see Figure 1). Eight out of 12 problems involve three premises and a conclusion, while the remaining four problems consist of four premises leading to a conclusion. All premises, as well as the conclusions resemble either biconditionals, exclusive disjunctions or conditionals. Two problems (4 and 6) include a redundant first premise. All premises are stated such that two subsequent statements contain one proposition in common, except of two problems (11 and 12), which are arranged in a non-sequential manner. In half of the problems, the conclusions logically follow from the premises, whereas in the other half, they do not. To avoid the influence of external knowledge and ensure content neutrality, Van der Henst et al. (2002) framed the problems around the presence of colored marbles in a box, with colors assigned randomly to each entity within a problem.
Language Models. We aim to investigate various factors that might impact the inferential strategies displayed by LLMs. These factors include the type of model, its size, and the emphasis on alignment during training Tunstall et al. (2023). Therefore, we assess a total of five models, consisting of three prominent open-access model types: Llama 2 Touvron et al. (2023) with model sizes of 7B, 13B, and 70B, the recently released Mistral-7B model Jiang et al. (2023), and Zephyr-7B Tunstall et al. (2023), an extension of Mistral-7B with a focus on intent alignment through fine-tuning with AI Feedback (AIF). For our evaluations, we utilize the publicly accessible model weights from the HuggingFace platform, specifically Llama-2-chat-hf333https://huggingface.co/meta-llama(7B, 13B, and 70B), Mistral-7B-Instruct-v0.2,444https://huggingface.co/mistralai/Mistral-7B-Instruct and zephyr-7b-beta.555https://huggingface.co/HuggingFaceH4/zephyr-7b-beta We consciously opt not to include proprietary models accessible via paid APIs, despite their reported superior performance in reasoning tasks Team et al. (2023). This methodological choice reflects our commitment to promoting transparent and reproducible scientific research. Note that in this work, we refer to the above models when using abbreviations such as LLaMA-2, Mistral-7B-Instruct or Zephyr-7B-.
Evaluation Setup. We prompt each model with a system message providing context about the task they are about to solve and the format in which they should answer (for the full prompt, see Figure 5 in the appendix). Analogous to Van der Henst et al. (2002), we inform the model of its participation in an experiment designed to explore reasoning processes, and instruct it to “think aloud” as it tackles the problem. In addition to the system message, we provide a user prompt that contains the problem description. In cases where the model does not accept system messages (such as Mistral-7B-Instruct-v0.2), we prepend the content of the system message to the user prompt. To prevent biasing the model towards a certain strategy, we refrain from providing few-shot examples, as done also by Leidinger et al. (2023). Instead, we elicit reasoning through zero-shot chain-of-thought prompting (“Let’s think step by step”) Kojima et al. (2023). Answers for each model are generated with nucleus sampling using Llama-2-chat-hf’s default values (, temperature ), as we found this configuration to work well for all models. To account for the statistical nature of language models, we ask each model to solve the set of propositional problems across 5 random seeds, resulting in a total of 60 responses per model.
We record all answers and manually evaluate them (a total of 300 responses) for strategies employed in their reasoning (see Figure 3 for an example). For each model response, we qualitatively evaluate for strategy and soundness. That is, we manually label the inferential strategies identified, and the logical validity of the model’s reasoning. In addition, we record whether the final answer is correct. In cases of faulty reasoning, we categorize the type of error. This comprehensive manual evaluation of model responses is independently conducted by two hired students with expertise in manual data annotation. To gauge the quality of the annotations, we report an overall Cohen’s Kappa value of . For details on the inter-annotator agreement of each label, we refer to Table 2 in the appendix. Further annotated examples can be found in Appendix C. Following the recommendations put forward by Leidinger et al. (2023), we make all input prompts, model responses, code and manual annotations publicly available upon acceptance of this paper.
Model |
[…]
green!20 Supposition Following |
[…]
yellow!20 Chain Construction |
[…]
orange!20 Compound Conclusion |
[…]
mylightblue Concatenation Strategy |
[…]
blue!20 Symbolic Strategy |
Correct Answer |
Sound Reasoning |
Zephyr-7B- |
Mistral-7B-Instruct |
LLaMA-2-7B |
LLaMA-2-13B |
LLaMA-2-70B |
Human Reasoner |
4 Results and Analysis
In this section, we present the results of our evaluation. We begin with a quantitative analysis of the inferential strategies employed by LLMs, as well as the logical validity of their reasoning. This is followed by a qualitative analysis providing a more in-depth examination of the models’ reasoning behavior.
4.1 Quantitative Analysis
Table 1 provides an overview of the frequencies with which large language models employ inferential strategies when navigating the problems of propositional logic described in Section 3. Our evaluation reveals that all models display strategies akin to those observed by Van der Henst et al. (2002). In particular, we find that, similar to humans, models commonly employ supposition following, chain construction and the compound strategy. In addition, we observe that models occasionally utilize the symbolic strategy, employing techniques from logical calculus to solve the tasks (see Section 2). Note that, similar to humans, models might switch from one strategy to another during a single problem, demonstrating multiple strategies within their responses (see Figure 19 in the appendix for an example). Surprisingly, we observe that distinct model families favor different inferential strategies. For instance, Zephyr-7B- predominantly employs supposition following, while Mistral-7B-Instruct is inclined towards drawing compound conclusions. In comparison, models from the Llama 2 series tend to rely on supposition following and chain construction, with negligible use of the compound strategy. Our analysis further reveals a discrepancy between the correctness of the models’ final answers and the logical soundness of their reasoning. While all models achieve an answer accuracy that approximately coincides with chance in our experimental setup, an analysis of their reasoning validity reveals a different picture: LLaMA-2-70B outperforms the other models by reasoning correctly in about 31.7% of cases, while Zephyr-7B- and Mistral-7B-Instruct produce sound reasoning in 25% of problems. We note that all models perform rather poorly on the propositional tasks, with LLaMA-2-7B failing entirely to construct sound arguments.
Human Reasoning. Van der Henst et al. (2002) compute the percentages with which human reasoners employ inferential strategies with respect to the total number of strategies observed in their experiment, and not with respect to the total number of problems considered. Thus, their reported values mainly reflect which strategies are favored more or less by the reasoner, but do not provide information about how frequently a strategy has been observed in the overall context. To make our findings comparable to the results of Van der Henst et al. (2002), we convert our results respectively (see values in parentheses in Table 1). We note that almost all models seem to favor supposition following to a higher degree than human reasoners, who employ this strategy in only about 21% of overall use. In contrast, humans seem to draw compound conclusions more readily, except for Mistral-7B-Instruct which shows a tendency more than twice as high. Overall, both LLMs and humans hardly employ the concatenation strategy. Interestingly, Van der Henst et al. (2002) report that all reasoners successfully solve the problems of propositional logic, though not always for the correct reasons. While the study does not provide data on the number of problems where humans reasoned correctly, the high success rate of human participants contrasts sharply with the performance of the models.
Effect of Model Size. Our evaluation of the Llama 2 series across three different model sizes—7B, 13B, and 70B parameters—demonstrates that model scale significantly influences the frequency with which strategies are employed by the model. In particular, we observe that with increasing model size, Llama 2 employs strategies more readily. Furthermore, larger models within the Llama 2 framework are observed to generate a greater number of sound reasoning traces. We interpret this trend as a result of the model’s improving proficiency in strategic reasoning as its scale increases.
Effect of Alignment. The alignment of a model’s response with human preferences is crucial to emulate human-like behavior Ouyang et al. (2022). Zephyr-7B- is an iteration of Mistral-7B that is fine-tuned with AI Feedback (AIF) for improved intent alignment Tunstall et al. (2023). In comparison to the observations made by Van der Henst et al. (2002), where besides the incremental diagram strategy (34%), chain construction was employed most frequently by humans, Zephyr-7B- demonstrates a marked preference for supposition following and significantly less engagement in chain construction. Moreover, it is noteworthy that among the evaluated models, Zephyr-7B- most frequently adopts the symbolic strategy, an approach not reported in human reasoners.
Sound Reasoning. As previously highlighted, the accuracy of a model’s final answer does not necessarily serve as a reliable indicator of its reasoning capability. In particular, we observe that models often arrive at correct answers, but through flawed reasoning processes (refer to Figure 10 in the appendix for an illustration). Interestingly, we also find instances where models provide incorrect final answers despite reasoning correctly (for an example, see Figure 16 in the appendix). Our analysis reveals only a moderate positive correlation between the accuracy of the models’ final answers and the logical soundness of their reasoning, with a Pearson correlation coefficient and a statistically significant p-value of less than . This observation underscores the need for more nuanced evaluation procedures, particularly in multiple-choice settings, where models might select the correct answer by chance rather than through rigorous reasoning.
In Figure 4, we explore the relationship between the inferential strategies employed by the models and the validity of their reasoning. For each strategy, we quantify the proportion of instances where the models’ reasoning is sound, compared to the overall application of that strategy. Our analysis reveals variability in the effectiveness with which different models apply various strategies. For example, Mistral-7B-Instruct tends to reason correctly when using approaches such as the chain, compound, or symbolic strategy, yet frequently encounters reasoning errors with supposition following. On the other hand, LLaMA-2-70B exhibits proficiency in supposition following, but struggles with the symbolic strategy.
4.2 Qualitative Analysis
We supplement our quantitative analysis by a more detailed qualitative analysis of the models’ reasoning behavior. Figure 3 depicts LLaMA-2-70B’s response to problem 1 of the task set. The response illustrates a frequently observed behavior. Initially, models tend to analyze the problem’s propositions, often by paraphrasing each premise and the conclusion to be evaluated. They then embark on a reasoning process, typically utilizing one of the previously mentioned strategies. In the example, LLaMA-2-70B employs chain construction, creating a logical chain of conditionals that leads from the antecedent of the final conclusion to its consequent, thereby correctly affirming the conclusion’s logical validity. A notable pitfall in such reasoning chains is the models’ occasional misinterpretation of logical negations, leading to erroneous chains like: A B; B C; therefore A C, where the negation in the first conditional is overlooked (for an illustrative case, refer to Figure 11 in the appendix). This behavior can be found across all models and aligns with previous work reporting difficulties of LLMs in understanding logical negations Truong et al. (2023).
When employing supposition following, models often fail to consider all implications of their assumptions. Instead, they tend to focus only on immediate inferences, while overlooking further consequences crucial for assessing the conclusion’s validity. This leads to models prematurely concluding the inability to definitively determine the logical validity of the final conclusion: “Based on our analysis, we cannot definitively say that the conclusion logically follows from the given statements” (see Figure 7 in the appendix for a respective example). Another source of error in supposition following involves models making improper suppositions, such as conjecturing about a marble not mentioned in the final conclusion, and deriving disjointed intermediate conclusions that do not aid in solving the problem. An example of this behavior can be found in Figure 8 in the appendix.
Finally, we identify two behaviors in models that mirror logical errors seen in human reasoners Van der Henst et al. (2002). First, models frequently attempt to prove an exclusive disjunction (A B) by only considering a single conditional case (A B), and second, they sometimes engage in the logical fallacy known as denial of the antecedent: A B; therefore A B (for illustrative examples, see Figures 12 and 13 in the appendix, respectively).
5 Related Work
Human Strategies in Deductive Reasoning. A considerable amount of research, especially within psychology and cognitive science, has explored how humans approach deductive reasoning tasks Schaeken et al. (2000). A prominent focus of these studies is on heuristics, which are cognitive shortcuts that individuals employ to arrive at satisfactory conclusions in deductive reasoning despite potential flaws in the underlying logic Kahneman et al. (1982); Evans (1989); Gigerenzer and Todd (1999); Davis (2018). For instance, Woodworth and Sells (1935) demonstrate that individuals tend to accept conclusions in syllogistic reasoning as valid when they share logical quantifiers with the premises, regardless of their actual logical validity. Nonetheless, such reliance on heuristics can result in errors and falls short of the level of strategic reasoning necessary to develop sound and coherent arguments Kahneman (2012). Further research has delved into more sophisticated strategies utilized by individuals in deductive reasoning. Based on the mental model theory Johnson-Laird (1986), Bucciarelli and Johnson-Laird (1999) identify a variety of strategies commonly employed by individuals in syllogistic reasoning. Byrne and Handley (1997) study strategies of individuals in knight-and-knave puzzles, where the truthfulness of statements made by hypothetical characters have to derived. Their experiments reveal that humans engage in both forward and backward inferences to navigate through potential solutions.
Human Reasoning Behavior in LLMs. Recent research has started to explore the extent to which LLMs mirror human-like reasoning behaviors. Dasgupta et al. (2023) demonstrate content-effects akin to those observed in human reasoning, where the deductive process is influenced by the content of the problem statement. Eisape et al. (2023) find that LLMs, similar to humans, exhibit biases such as ordering effects in syllogistic reasoning tasks. Several other studies have delved into the prevalence of biases and heuristics within LLMs Binz and Schulz (2023); Talboy and Fuller (2023); Shaki et al. (2023); Suri et al. (2023). However, to the best of our knowledge, we are the first who study the presence of more sophisticated human strategies in the context of LLM-based deductive reasoning.
6 Conclusion
In this paper, we examine the inferential strategies employed by LLMs in solving problems of propositional logic. Through a comprehensive evaluation of their reasoning behavior, we demonstrate that LLMs adopt strategies akin to those observed in human reasoners. Our quantitative analysis reveals that the frequency with which a model adopts a specific strategy strongly depends on its type, size, and fine-tuning procedure. Moreover, our analysis suggests that the accuracy of a model’s final conclusions does not adequately capture its reasoning capabilities, underscoring the importance of a more sophisticated evaluation framework that includes the model’s reasoning paths. We also provide a qualitative analysis of typical reasoning behaviors among models, pinpointing prevalent errors such as difficulties in understanding negations or recognizing all implications of a supposition.
7 Limitations
While our work contributes to the understanding of reasoning processes in large language models by demonstrating that these models, similar to humans, employ inferential strategies in propositional logic, it encompasses several limitations that could be addressed in future work.
Task setup. Our study is constrained by a limited set of problems, designed within a fixed framework that revolves around hypothesis validation based on 3-4 statements of propositional logic. We employ a constant and neutral content, disregarding potential content-effects on the models’ reasoning behavior, as shown by Dasgupta et al. (2023). Similarly, we have not yet examined factors such as the complexity of the problems, the differences between hypothesis validation and generation, and the impact of logical connectives utilized in the premises. We belief that these factors are worth investigating and leave a detailed examination to future work.
Evaluation Framework. The extent of our manual evaluation is limited by both the number of samples reviewed and the quantity of annotators involved. Despite our efforts to maximize the use of available resources, these constraints may affect the scalability and reliability of our results. Additionally, we instruct all models through zero-shot chain-of-thought prompting (“Let’s think step by step”) Kojima et al. (2023). Exploring alternative reasoning frameworks, such as Tree of Thoughts Yao et al. (2023) or Graph of Thoughts Besta et al. (2023), could provide valuable insights into their influence on model behavior and the inferential strategies adopted. Based on our annotated data, we endeavored to develop a classifier capable of automatically identifying the inferential strategies employed in the models’ output, which was intended to complement our manual evaluation setup. However, due to the complexity of the task and limited size of our annotated dataset, our classifier struggled with generalization to new, unseen responses. In future endeavors, we aim to allocate more resources towards expanding our manual annotation efforts and explore this direction further. Finally, our study predominantly offers a behavioral analysis and does not delve into the mechanistic aspects that might explain the diversity in strategy usage by the models. Investigating how attention patterns of models with respect to specific logical connectives might influence their choice of reasoning strategy presents a compelling direction for future research.
Acknowledgements
We would like to thank the MaiNLP lab members for their insightful feedback, specifically Diego Frassinelli, Rob van der Goot, Siyao Peng, Robert Litschko, Jian Lan, Daniela Teodorescu, Xinpeng Wang, Verena Blaschke, Elena Senger, Elisa Bassignana, and Max Müller-Eberstein. Furthermore, we would like to express our gratitude to Huangyan Shan for her valuable work and support in data annotation. Our appreciation extends to the anonymous reviewers for their comments and suggestions. Lastly, we acknowledge the support for BP through the ERC Consolidator Grant DIALECT 101043235.
Appendix A Additional Experimental Details
In this section, we provide additional details about the experimental setup, including supplementary information about the problem formulations and prompts utilized.
[…]
green!20 Supposition Following |
[…]
yellow!20 Chain Construction |
[…]
orange!20 Compound Conclusion |
[…]
mylightblue Concatenation Strategy |
[…]
blue!20 Symbolic Strategy |
Correct Answer |
Sound Reasoning |
Zephyr-7B- |
Mistral-7B-Instruct |
LLaMA-2-7B |
LLaMA-2-13B |
LLaMA-2-70B |
A.1 Task Prompts
Figure 5 displays the task prompt and problem formulations employed in assessing the language models described in Section 3. Note that the prompt template, i.e. special tokens and their arrangements, might vary depending on the specific language model used. Within the task prompt (provided in the upper box), the problem statements and conclusion for a given problem are replaced with the corresponding problem formulations found in the lower gray boxes. In the final version of the prompt, the phrase “colorA iff colorB” is expanded to “There is a colorA marble in the box if and only if there is a colorB marble in the box”. Similarly, “colorA xor colorB” is interpreted as “There is either a colorA marble or a colorB marble in the box, but not both”, and “If colorA then colorB” is articulated as “If there is a colorA marble in the box, then there is a colorB marble in the box”.
A.2 Annotator Instructions
Our assessment of model responses involves a comprehensive independent review by two students who are specialized in the field of natural language processing and have expertise in manual data annotation. To ensure a high quality of annotations, we offer comprehensive training to both annotators. This training includes detailed explanations and extensive examples of the strategies identified by Van der Henst et al. (2002), complemented by a session dedicated to clarifying any questions that may emerge. Subsequently, the annotators are tasked with independently annotating practice examples, which serves to highlight and address any ambiguities in the annotation process. Only when both annotators are confident in their understanding of each strategy do we proceed. We instruct both annotators to independently go through each model response and mark parts where they identify a certain strategy to be employed. Each strategy is marked in a unique color code, which is afterwards converted into labels that signify the use of a particular strategy. In addition, we instruct both annotators to label whether the reasoning is sound, and the final conclusion of the model is correct. Furthermore, we ask them to classify any logical errors identified within the reasoning process. To maintain a high standard of annotation quality, annotators are instructed to review the model responses twice.
A.3 Inter-Annotator Agreement
To assess the reliability of our manual evaluation process (see Section 3), we quantify the inter-annotator agreement by calculating Cohen’s Kappa for each category and model, as illustrated in Table 2. Generally, the results indicate an almost perfect level of agreement across all categories and models, with Cohen’s Kappa values ranging from . An exception is observed in the case of the concatenation strategy applied by LLaMA-2-7B, for which we report a substantial agreement level, with a Kappa value of , slightly below the threshold for almost perfect agreement.
A.4 Model Details
We report further details about the models used in this study in Table 4. In particular, we provide information about the number of parameters, context length and fine-tuning procedure for each model.
Appendix B Additional Quantitative Results
In this segment, we present supplementary findings from our quantitative evaluation. Table 3 illustrates the frequencies with which the different language models employ inferential strategies when navigating the problems of propositional logic, as outlined in Section 3. Values denote percentages averaged across five distinct random seeds, accompanied by their standard deviation. Furthermore, we detail the proportions of correct final conclusions and sound reasoning. Note that all percentages are calculated relative to the overall count of tasks within the experimental framework.
Appendix C Annotated Model Responses
Within this section, we showcase examples of model responses that exemplify each inferential strategy identified in our study, as depicted in figures 6-19. Each figure is organized with the problem statement at the top, the model’s response on the lower left, and the annotators’ comments to the lower right. For an extensive array of model responses and annotations, we invite readers to explore our data repository, which we publish upon acceptance of this paper.
C.1 Supposition Following
Figures from 6 to 8 demonstrate the application of supposition following by various models. For instance, Figure 6 presents LLaMA-2-70B’s approach to problem 7, where the model supposes the absence of a blue marble in the box and logically infers the implications of this assumption to reach the valid conclusion. On the other hand, Figure 7 depicts Mistral-7B-Instruct’s response to the same problem, where the model considers various combinations of marble in the box, drawing immediate conclusions that follow from the premises at hand. However, it does not explore deeper ramifications of these suppositions, thereby failing to deduce the validity of the conclusion. This showcases a common behavior we observe in models that employ supposition following unsuccessfully. In Figure 8 the model approaches problem 9 by assuming the presence of an olive marble in the box, yet inferring disjointed intermediate conclusions that do not aid in solving the problem, thus failing to prove the logical validity of the problem.
Model |
[…]
green!20 Supposition Following |
[…]
yellow!20 Chain Construction |
[…]
orange!20 Compound Conclusion |
[…]
mylightblue Concatenation Strategy |
[…]
blue!20 Symbolic Strategy |
Correct Answer |
Sound Reasoning |
Zephyr-7B- |
Mistral-7B-Instruct |
LLaMA-2-7B |
LLaMA-2-13B |
LLaMA-2-70B |
C.2 Chain Construction
Figures 9 to 13 illustrate instances where models employ chain construction to navigate the problems of propositional logic. In Figure 9, LLaMA-2-70B adeptly forms a chain of conditional statements that bridge the antecedent of the conclusion to its consequent, effectively validating the conclusion’s logical soundness. Conversely, Figure 10 depicts a logical chain in which LLaMA-2-70B erroneously concludes the nonexistence of a white marble based on the absence of a red marble, despite an exclusive disjunction linking the two. Despite this logical misstep, the model’s final conclusion remains accurate, highlighting the discrepancy between the model’s final answer and the soundness of its reasoning. In Figure 11, LLaMA-2-13B constructs a chain correctly linking the antecedent of the final conclusion to its consequent. Nonetheless, it overlooks the negation present in one of the conditionals, resulting in a compromised reasoning chain. Figure 12 presents a scenario where the model incorrectly attempts to validate an exclusive disjunction solely through a singular conditional sequence, a reasoning error not uncommon among human reasoners Van der Henst et al. (2002). Lastly, Figure 13 highlights LLaMA-2-70B’s engagement in the inverse fallacy, inferring W G from the conditional W G, mirroring a logical misjudgment frequently observed in human reasoning processes.
C.3 Compound Strategy
The compound strategy is illustrated in Figures 14 to 16. Figure 14 presents Mistral-7B-Instruct’s approach to problem 9, where it infers a biconditional relationship between the purple and olive marble from the first two premises. On the other hand, Figure 15 shows LLaMA-2-70B’s response to the same problem, formulating a sequence of compound inferences beyond the initial biconditional deduction, culminating in the correct final answer. Additionally, Figure 16 illustrates Mistral-7B-Instruct’s approach to problem 8, in which the model initially generates compound conclusions derived from the problem statements, followed by supposition following to explore the implications that the absence of an olive marble might have. However, despite the model’s sound reasoning, its final answer is incorrect.
C.4 Concatenation Strategy
Figure 17 demonstrates the concatenation strategy, where Mistral-7B-Instruct concatenates two intermediate deductions to form a single statement. It then uses the concatenated statement to infer the invalidity of the conclusion.
C.5 Symbolic Strategy
The symbolic strategy is exemplified in Figure 18, where LLaMA-2-70B employs a truth table to assess the conclusion’s validity, albeit with errors leading to an incorrect result. Conversely, Figure 19 shows Mistral-7B-Instruct’s application of chain construction followed by the symbolic strategy. The model makes false inferences while employing chain construction, and further errs in its validation through logical calculus.
Model |
Base Model |
Parameters |
Context Length |
Tokens |
Fine-tuning |
Zephyr-7B- |
Mistral |
7B |
8192 tokens |
dSFT, AIF |
Mistral-7B-Instruct |
Mistral |
7B |
8192 tokens |
SFT |
LLaMA-2-7B-Chat |
LLaMA-2 |
7B |
4K tokens |
2.0T |
SFT, RLHF |
LLaMA-2-13B-Chat |
LLaMA-2 |
13B |
4K tokens |
2.0T |
SFT, RLHF |
LLaMA-2-70B-Chat |
LLaMA-2 |
70B |
4K tokens |
2.0T |
SFT, RLHF |
Response:
Sure, I’d be happy to help! Here’s my reasoning process for this problem:
1. First, I’ll start by analyzing the given statements:
Statement 1: "There is a blue marble in the box if and only if there is a red marble in the box"
This statement tells me that the presence of a blue marble in the box is dependent on the presence of a red marble in the box. If there’s no red marble, there can’t be a blue marble.
Statement 2: "Either there is a red marble in the box or else there is a white marble in the box, but not both"
This statement tells me that the presence of a red marble and a white marble in the box are mutually exclusive. If there’s a red marble, there can’t be a white marble, and vice versa.
Statement 3: "There is a white marble in the box if and only if there is a pink marble in the box"
This statement tells me that the presence of a white marble in the box is dependent on the presence of a pink marble in the box. If there’s no pink marble, there can’t be a white marble.
2. Now, let’s consider the conclusion: "If there is not a blue marble in the box then there is a pink marble in the box"
Response:
To determine if the conclusion logically follows from the given statements, let’s first understand the given statements and the logical relationships between them:
1. Statement 1: "There is a blue marble in the box if and only if there is a red marble in the box."
- This statement is equivalent to saying that the presence of a blue marble and the presence of a red marble are mutually exclusive and exhaustive.
2. Statement 2: "Either there is a red marble in the box or else there is a white marble in the box, but not both."
- This statement tells us that there cannot be both a red and a white marble in the box.
3. Statement 3: "There is a white marble in the box if and only if there is a pink marble in the box."
- This statement is equivalent to saying that the presence of a white marble and the presence of a pink marble are mutually exclusive and exhaustive.
Now let’s reason about the conclusion:
1. Conclusion: "If there is not a blue marble in the box then there is a pink marble in the box."
Response:
Based on the given statements, here’s my reasoning:
1. First, let’s analyze the given statements one by one:
a) Statement 1: "There is a purple marble in the box if and only if there is a blue marble in the box"
b) Statement 2: "There is a blue marble in the box if and only if there is an olive marble in the box"
c) Statement 3: "Either there is an olive marble in the box or else there is a red marble in the box, but not both"
d) Statement 4: "Either there is a red marble in the box or else there is a green marble in the box, but not both"
2. Since we are trying to prove that "If there is a purple marble in the box then there is a green marble in the box", let’s try to find a way to connect purple marbles with green marbles using the given statements.
3. Unfortunately, there doesn’t seem to be a direct connection between purple marbles and green marbles in the given statements.
4. However, let’s see if we can find any indirect connections. Based on Statement 1, we know that there is a purple marble in the box if and only if there is a blue marble in the box.
5. Now, let’s look at Statement 2. We know that there is a blue marble in the box if and only if there is an olive marble in the box.
6. From Statement 3, we know that there cannot be both an olive marble and a red marble in the box. However, Statement 4 tells us that there cannot be both a red marble and a green marble in the box either.
Response:
Sure, I’d be happy to help! Here’s my reasoning process for this problem:
1. First, I’ll start by analyzing the given statements:
Statement 1: "There is a purple marble in the box if and only if there is a blue marble in the box"
This statement tells me that the presence of a purple marble in the box is dependent on the presence of a blue marble in the box. If there is a blue marble, then there must be a purple marble, and vice versa.
Statement 2: "There is a blue marble in the box if and only if there is an olive marble in the box"
This statement tells me that the presence of a blue marble in the box is dependent on the presence of an olive marble in the box. If there is an olive marble, then there must be a blue marble, and vice versa.
Statement 3: "Either there is an olive marble in the box or else there is a red marble in the box, but not both"
This statement tells me that the presence of an olive marble and the presence of a red marble in the box are mutually exclusive. If there is an olive marble, then there cannot be a red marble, and vice versa.
Statement 4: "Either there is a red marble in the box or else there is a green marble in the box, but not both"
This statement tells me that the presence of a red marble and the presence of a green marble in the box are also mutually exclusive. If there is a red marble, then there cannot be a green marble, and vice versa.
2. Now, let’s consider the conclusion: "If there is a purple marble in the box then there is a green marble in the box"
My reasoning process for this conclusion is as follows:
Response:
Sure! Let’s go through the statements and see if we can determine whether the conclusion logically follows.
1. "There is a blue marble in the box if and only if there is a red marble in the box"
This statement tells us that the presence of a blue marble in the box is dependent on the presence of a red marble. If there is a red marble, then there must be a blue marble, and if there is no red marble, then there cannot be a blue marble.
2. "Either there is a red marble in the box or else there is a white marble in the box, but not both"
This statement tells us that the presence of a red marble and a white marble in the box are mutually exclusive. If there is a red marble, then there cannot be a white marble, and if there is a white marble, then there cannot be a red marble.
3. "There is a white marble in the box if and only if there is a pink marble in the box"
This statement tells us that the presence of a white marble in the box is dependent on the presence of a pink marble. If there is a pink marble, then there must be a white marble, and if there is no pink marble, then there cannot be a white marble.
Now, let’s consider the conclusion: "If there is not a blue marble in the box then there is a pink marble in the box"
Response:
Sure, I’d be happy to help with this experiment! Here’s my reasoning process for each statement and the conclusion:
1. There is a brown marble in the box if and only if there is a yellow marble in the box
This statement means that if there is a brown marble in the box, then there must also be a yellow marble in the box. Conversely, if there is a yellow marble in the box, then there must also be a brown marble in the box.
2. Either there is a yellow marble in the box or else there is a green marble in the box, but not both
This statement means that if there is a yellow marble in the box, then there cannot also be a green marble in the box. Similarly, if there is a green marble in the box, then there cannot also be a yellow marble in the box.
3. There is a green marble in the box if and only if there is a purple marble in the box
This statement means that if there is a green marble in the box, then there must also be a purple marble in the box. Conversely, if there is a purple marble in the box, then there must also be a green marble in the box.
4. There is a purple marble in the box if and only if there is an olive marble in the box
This statement means that if there is a purple marble in the box, then there must also be an olive marble in the box. Conversely, if there is an olive marble in the box, then there must also be a purple marble in the box.
Now, let’s consider the conclusion:
Conclusion: If there is a brown marble in the box then there is an olive marble in the box
Based on the statements we’ve been given, we can logically deduce the conclusion as follows:
Response:
Sure! Let’s break this down step by step.
1. "There is a purple marble in the box if and only if there is a yellow marble in the box"
This statement implies that the presence of a purple marble in the box is dependent on the presence of a yellow marble in the box. If there is no yellow marble, there cannot be a purple marble.
2. "There is a yellow marble in the box if and only if there is a blue marble in the box"
This statement implies that the presence of a yellow marble in the box is dependent on the presence of a blue marble in the box. If there is no blue marble, there cannot be a yellow marble.
3. "Either there is a blue marble in the box or else there is an orange marble in the box, but not both"
This statement implies that the presence of a blue marble and an orange marble in the box are mutually exclusive. If there is a blue marble, there cannot be an orange marble, and vice versa.
Now, let’s consider the conclusion: "Either there is a purple marble in the box or else there is an orange marble in the box, but not both"
Response:
Sure! Let’s break down the problem and think through each step.
1. First, we have the statement "There is a blue marble in the box if and only if there is a brown marble in the box." This means that if there is a blue marble in the box, then there must be a brown marble in the box, and vice versa.
2. Next, we have the statement "There is a white marble in the box if and only if there is a green marble in the box." This means that if there is a white marble in the box, then there must be a green marble in the box, and vice versa.
3. The third statement is "Either there is a brown marble in the box or else there is a white marble in the box, but not both." This means that if there is a brown marble in the box, then there cannot be a white marble in the box, and vice versa.
4. The fourth statement is "There is a green marble in the box if and only if there is a purple marble in the box." This means that if there is a green marble in the box, then there must be a purple marble in the box, and vice versa.
5. Now, let’s consider the conclusion: "If there is a blue marble in the box then there is a purple marble in the box."
Let’s think about how we can use the given statements to determine whether this conclusion is true or false.
Response:
First, let’s analyze the given statements:
1. Statement 1: There is a purple marble in the box if and only if there is a blue marble in the box
2. Statement 2: There is a blue marble in the box if and only if there is an olive marble in the box
3. Statement 3: Either there is an olive marble in the box or else there is a red marble in the box, but not both
4. Statement 4: Either there is a red marble in the box or else there is a green marble in the box, but not both
Let’s examine the relationship between the statements:
Response:
Sure! Let’s break it down step by step.
1. "There is a purple marble in the box if and only if there is a blue marble in the box"
This statement implies that the presence of a purple marble in the box is dependent on the presence of a blue marble in the box. If there is no blue marble, there cannot be a purple marble.
2. "There is a blue marble in the box if and only if there is an olive marble in the box"
This statement implies that the presence of a blue marble in the box is dependent on the presence of an olive marble in the box. If there is no olive marble, there cannot be a blue marble.
3. "Either there is an olive marble in the box or else there is a red marble in the box, but not both"
This statement implies that the presence of an olive marble and a red marble in the box are mutually exclusive. If there is an olive marble, there cannot be a red marble, and vice versa.
4. "Either there is a red marble in the box or else there is a green marble in the box, but not both"
This statement implies that the presence of a red marble and a green marble in the box are also mutually exclusive. If there is a red marble, there cannot be a green marble, and vice versa.
Now, let’s evaluate the conclusion:
Conclusion: If there is a purple marble in the box then there is a green marble in the box
Response:
To determine if the conclusion logically follows from the given statements, we need to use the principles of logical reasoning. Let’s analyze the statements one by one and then consider the conclusion.
1. Statement 1: Either there is an olive marble in the box or else there is a brown marble in the box, but not both
2. Statement 2: There is a brown marble in the box if and only if there is a gray marble in the box
3. Statement 3: Either there is a gray marble in the box or else there is a maroon marble in the box, but not both
First, let’s understand the implications of Statement 2. It states that if there is a brown marble in the box, then there is a gray marble in the box, and conversely, if there is a gray marble in the box, then there is a brown marble in the box. This is an if and only if statement, meaning both conditions are equivalent.
Response:
Based on the given statements, here is my reasoning:
1. The first statement tells us that there is a gray marble in the box if and only if there is a yellow marble in the box.
2. The second statement tells us that either there is a yellow marble in the box or there is an olive marble in the box, but not both.
3. The third statement tells us that there is an olive marble in the box if and only if there is a black marble in the box.
4. From the second statement, we know that if there is a yellow marble in the box, then there cannot be an olive marble in the box.
5. From the third statement, we know that if there is an olive marble in the box, then there is a black marble in the box.
Response:
Sure, I’d be happy to help! Here’s my reasoning process for this problem:
1. First, I’ll start by analyzing the given statements:
- Statement 1: "There is a green marble in the box if and only if there is a purple marble in the box"
- This statement establishes a direct relationship between the presence of a green marble and a purple marble in the box.
- Statement 2: "If there is a purple marble in the box then there is a gray marble in the box"
- This statement establishes a conditional relationship between the presence of a purple marble and a gray marble in the box.
- Statement 3: "Either there is a gray marble in the box or else there is a yellow marble in the box, but not both"
- This statement establishes an exclusive OR relationship between the presence of a gray marble and a yellow marble in the box.
2. Now, let’s consider the conclusion: "Either there is a green marble in the box or else there is a yellow marble in the box, but not both"
Response:
First, we need to understand the relationships between the marbles and the statements. Here’s a summary:
- There is a gray marble in the box if and only if there is a yellow marble in the box
- This statement creates a connection between the gray marble and the yellow marble. If one is present, the other is as well.
- Either there is a yellow marble in the box or else there is an olive marble in the box, but not both
- This statement introduces two possible scenarios. Either the yellow marble is present, or the olive marble is present, but not both.
- There is an olive marble in the box if and only if there is a black marble in the box
- This statement creates another connection, but this time between the olive marble and the black marble. If one is present, the other is as well.
Now, let’s see if the conclusion follows logically from these statements.
- If there is a yellow marble in the box then there is a black marble in the box

### v2

@@
 The inferential strategy employed is significantly influenced by the model family, as different families favor different approaches.
-Models are often right but for the wrong reasons: the accuracy of a model, that is the number of correct final answers, does not reflect whether its reasoning is sound, i.e.
+Models are often right but for the wrong reasons: the accuracy of a model, that is the number of correct final conclusions, does not reflect whether its reasoning is sound, i.e.
 logically follows from the statements at hand.
 The strategy employed by a model is closely related to the soundness of its reasoning, where certain strategies lead to correct reasoning and others tend to introduce errors.
-In contrast to human reasoners, models occasionally adopt a symbolic strategy, where formal logical calculus is employed to solve the problems of propositional logic.
+In contrast to humans, models occasionally adopt a symbolic strategy, where formal logical calculus is employed to solve the propositional logic problem at hand.
 Through this work, we hope to advance the understanding of reasoning in LLMs.
@@
 Van der Henst et al.
-(2002) discovered five strategies reasoners commonly utilize to navigate the problems, offering insights into their inferential mechanisms employed in propositional reasoning.
+(2002) discovered five strategies reasoners commonly utilize to navigate the problems, offering insights into their inferential mechanisms employed during propositional reasoning.
 In the following, we give a short description of each strategy (illustrated in Figure 2).
@@
 Incremental Diagram.
-This strategy involves the creation of a comprehensive diagram that keeps track of all potential outcomes compatible with the statements of the problem.
+This strategy involves the creation of a comprehensive diagram that keeps track of all potential outcomes compatible with the premises of the problem.
 During the reasoning process, individuals progressively increment their diagrams to incorporate new information derived (see left box in Figure 2).
 The result is a single diagram that records a variety of possibilities compatible with the premises, often including even those that might be irrelevant to the task.111In contrast to Van der Henst et al.
-(2002), we in fact observe no single occurrence of the incremental diagram strategy in LLMs, despite the authors finding this strategy being employed most frequently by humans.
-We belief that this discrepancy stems from the use of pen and paper in human assessments, implicitly encouraging diagrammatic reasoning.
+(2002), we in fact observe no single occurrence of the incremental diagram strategy in LLMs, despite the authors finding that this strategy is most frequently employed by humans.
+We believe that this discrepancy stems from the use of pen and paper in human assessments, implicitly encouraging diagrammatic reasoning.
+Exploring how this observation changes with vision-language models would be an intriguing area for future research.
 Supposition Following.
@@
 by assuming a marble of a certain color.
-They then trace the consequences of that supposition following from the statements at hand, as illustrated in the second box from the left of Figure 2.
+Subsequently, they trace the implications of that supposition, logically following from the premises at hand, as illustrated in the upper second box from the left of Figure 2.
 The result is a sequence of literals (in this case, marbles of a certain color) without logical connectives.
-The success of supposition following strongly depends on the supposition made by the reasoner.
-Some suppositions might lead to inferences that are relevant to the problem, while others might lead to irrelevant conclusions.
+The efficiency and success of supposition following strongly depends on the supposition made by the reasoner.
+While some suppositions lead to inferences that are relevant to the problem, others might lead to irrelevant conclusions.
 Chain Construction.
-When employing this strategy, reasoners construct a chain of conditional statements derived from the premises at hand or from intermediate deductions.
-An example of chain construction is displayed in the central box of Figure 2.
+When employing this strategy, reasoners construct a chain of conditional statements derived either from the premises in the problem description or from intermediate deductions.
+An example of chain construction is displayed in the lower second box from the left of Figure 2.
 Premises are converted into a chain of conditional statements that are linked by their entities.
@@
 This process yields a series of novel conclusions, each building upon the preceding ones.
-An example illustrating this strategy is given in the upper second box from the right of Figure 2.
-Based on the first two premises, the compound conclusion: “If red then not blue.” is inferred, and then used to draw another compound conclusion (“If red or pink then not blue.”) together with the last premise of the problem statement.
+An illustrative example of this strategy is given in the upper second box from the right of Figure 2.
+Based on the first two premises, the compound conclusion: “If blue then not red.” is inferred, and then used to draw another compound conclusion (“If blue then not red and not pink.”) together with the last premise of the problem statement.
 Concatenation Strategy.
@@
 3 Experimental Setup
+Statements: 1.
+Either there is a white marble in the box or else there is a black marble in the box, but not both.
+2.
+Either there is a black marble in the box or else there is a pink marble in the box, but not both.
+3.
+There is a pink marble in the box if and only if there is a gray marble in the box.
+Conclusion: If there is a white marble in the box then there is a gray marble in the box.
+Response:
+Sure, I’d be happy to help! Here’s my reasoning process for this problem:
+1. First, I’ll start by analyzing the given statements:
+From Statement 1, we know that the box contains either a white marble or a black marble, but not both.
+From Statement 2, we know that the box contains either a black marble or a pink marble, but not both.
+From Statement 3, we know that if there is a pink marble in the box, then there must be a gray marble in the box.
+2. Now, let’s consider the conclusion: If there is a white marble in the box then there is a gray marble in the box.
+Can we logically deduce this conclusion from the given statements?
 Task Overview.
@@
 (2002) (an overview of each problem can be found in Figure 5 in the appendix).
-In each problem, models are presented with a set of statements (or premises) and must determine whether a given conclusion logically follows (for an example, see Figure 1).
+For each problem, models are presented with a set of statements (or premises) and must determine whether a given conclusion logically follows (for an example, see Figure 1).
 Eight out of 12 problems involve three premises and a conclusion, while the remaining four problems consist of four premises leading to a conclusion.
@@
 All premises are stated such that two subsequent statements contain one proposition in common, except of two problems (11 and 12), which are arranged in a non-sequential manner.
-In half of the problems, the conclusions logically follow from the premises, whereas in the other half, they do not.
+For half of the problems, the conclusions logically follow from the premises, whereas for the other half, they do not.
 To avoid the influence of external knowledge and ensure content neutrality, Van der Henst et al.
@@
 To account for the statistical nature of language models, we ask each model to solve the set of propositional problems across 5 random seeds, resulting in a total of 60 responses per model.
+Our code is publicly available at: https://github.com/mainlp/inferential-strategies.
 We record all answers and manually evaluate them (a total of 300 responses) for strategies employed in their reasoning (see Figure 3 for an example).
@@
 Following the recommendations put forward by Leidinger et al.
-(2023), we make all input prompts, model responses, code and manual annotations publicly available upon acceptance of this paper.
-Model |
-[…]
-green!20 Supposition Following |
-[…]
-yellow!20 Chain Construction |
-[…]
-orange!20 Compound Conclusion |
-[…]
-mylightblue Concatenation Strategy |
-[…]
-blue!20 Symbolic Strategy |
-Correct Answer |
-Sound Reasoning |
-Zephyr-7B- |
-Mistral-7B-Instruct |
-LLaMA-2-7B |
-LLaMA-2-13B |
-LLaMA-2-70B |
-Human Reasoner |
+(2023), we make all input prompts, model responses and manual annotations publicly available at: huggingface.co/datasets/mainlp/inferential_strategies.
+| Model | Supposition Following | Chain Construction | Compound Conclusion | Concatenation Strategy | Symbolic Strategy | Correct Answer | Sound Reasoning |
+| Zephyr-7B- | |||||||
+| Mistral-7B-Instruct | |||||||
+| LLaMA-2-7B | |||||||
+| LLaMA-2-13B | |||||||
+| LLaMA-2-70B | |||||||
+| Human Reasoner† |
 4 Results and Analysis
@@
 We begin with a quantitative analysis of the inferential strategies employed by LLMs, as well as the logical validity of their reasoning.
-This is followed by a qualitative analysis providing a more in-depth examination of the models’ reasoning behavior.
+This is followed by a qualitative analysis providing a more in-depth examination of the models’ reasoning.
 4.1 Quantitative Analysis
@@
 Surprisingly, we observe that distinct model families favor different inferential strategies.
-For instance, Zephyr-7B- predominantly employs supposition following, while Mistral-7B-Instruct is inclined towards drawing compound conclusions.
-In comparison, models from the Llama 2 series tend to rely on supposition following and chain construction, with negligible use of the compound strategy.
+For instance, Zephyr-7B- predominantly employs supposition following, while Mistral-7B-Instruct is equally inclined towards drawing compound conclusions.
+In contrast, models from the Llama 2 series tend to rely on supposition following and chain construction, with negligible use of the compound strategy.
 Our analysis further reveals a discrepancy between the correctness of the models’ final answers and the logical soundness of their reasoning.
-While all models achieve an answer accuracy that approximately coincides with chance in our experimental setup, an analysis of their reasoning validity reveals a different picture: LLaMA-2-70B outperforms the other models by reasoning correctly in about 31.7% of cases, while Zephyr-7B- and Mistral-7B-Instruct produce sound reasoning in 25% of problems.
+While all models achieve an answer accuracy that approximately coincides with chance in our experimental setup, an analysis of their reasoning validity reveals a different picture: LLaMA-2-70B outperforms the other models by reasoning correctly in about 31.7% of cases, while Zephyr-7B- and Mistral-7B-Instruct produce sound reasoning in 25% of the problems.
 We note that all models perform rather poorly on the propositional tasks, with LLaMA-2-7B failing entirely to construct sound arguments.
@@
 Our analysis reveals only a moderate positive correlation between the accuracy of the models’ final answers and the logical soundness of their reasoning, with a Pearson correlation coefficient and a statistically significant p-value of less than .
-This observation underscores the need for more nuanced evaluation procedures, particularly in multiple-choice settings, where models might select the correct answer by chance rather than through rigorous reasoning.
+This observation aligns with findings from previous studies Ye and Durrett (2022); Creswell and Shanahan (2022) and underscores the need for more nuanced evaluation procedures, particularly in multiple-choice settings, where models might select the correct answer by chance rather than through rigorous reasoning.
 In Figure 4, we explore the relationship between the inferential strategies employed by the models and the validity of their reasoning.
@@
 However, to the best of our knowledge, we are the first who study the presence of more sophisticated human strategies in the context of LLM-based deductive reasoning.
+Faithful Reasoning with LLMs.
+Large language models can be instructed to explain the reasoning process by which they derive their final conclusions Wei et al.
+(2022); Kojima et al.
+(2022).
+However, several studies indicate that these self-explanations might not always be faithful, i.e.
+accurately represent the model’s underlying reasoning process Jacovi and Goldberg (2020); Agarwal et al.
+(2024); Lyu et al.
+(2024).
+For instance, Turpin et al.
+(2023) demonstrate that LLMs such as GPT-3.5 OpenAI (2023) and Claude 1 Anthropic (2023) often fail to mention biasing features in their input that significantly influence their decisions.
+Instead, the models produce plausible yet misleading explanations that give a false account of the underlying decision process.
+Lanham et al.
+(2023) probe the faithfulness of explanations by evaluating how the final conclusions of LLMs change when rationales are truncated or errors are introduced.
+Their findings reveal that the extent to which models rely on their rationales varies strongly across models and tasks.
+Matton et al.
+(2024) propose a method to quantify the faithfulness of explanations based on high-level concepts in the models’ input that influence decision-making.
+By measuring the difference between the set of concepts that LLMs deem influential and the set that truly are, instances of unfaithfulness could be identified, including cases where LLMs overlook the impact of social biases in their decision-making processes.
+Another important consideration is whether the model’s final conclusion aligns with its preceding explanation.
+As highlighted in Section 4.1, a correct conclusion might not always be the product of a logically sound reasoning trace, particularly in multiple-choice setups.
+Conversely, a sound rationale may not always lead to a logically consistent answer.
+Related work by Ye and Durrett (2022) indicates that in question-answering and natural language inference tasks, explanations generated by LLMs such as OPT Zhang et al.
+(2022) and GPT-3 Brown et al.
+(2020) often do not entail the models’ final conclusions.
+Further studies aim to enhance the models’ faithfulness, for instance by enforcing causality from proof generation to entailment prediction.
+This can be achieved by either restricting the model’s context Sanyal et al.
+(2022); Creswell and Shanahan (2022); Radhakrishnan et al.
+(2023), or by utilizing deterministic tools that are inherently faithful by design Lyu et al.
+(2023).
 6 Conclusion
@@
 7 Limitations
-While our work contributes to the understanding of reasoning processes in large language models by demonstrating that these models, similar to humans, employ inferential strategies in propositional logic, it encompasses several limitations that could be addressed in future work.
+While our work contributes to the understanding of reasoning processes in large language models by demonstrating that these models employ inferential strategies in propositional logic similar to humans, it encompasses several limitations that could be addressed in future work.
 Task setup.
@@
 Similarly, we have not yet examined factors such as the complexity of the problems, the differences between hypothesis validation and generation, and the impact of logical connectives utilized in the premises.
-We belief that these factors are worth investigating and leave a detailed examination to future work.
+We believe that these factors are worth investigating and leave a detailed examination to future work.
 Evaluation Framework.
@@
 Finally, our study predominantly offers a behavioral analysis and does not delve into the mechanistic aspects that might explain the diversity in strategy usage by the models.
-Investigating how attention patterns of models with respect to specific logical connectives might influence their choice of reasoning strategy presents a compelling direction for future research.
+Investigating how model-internal mechanisms might influence their choice of reasoning strategy presents a compelling direction for future research.
 Acknowledgements
@@
 In this section, we provide additional details about the experimental setup, including supplementary information about the problem formulations and prompts utilized.
-[…]
-green!20 Supposition Following |
-[…]
-yellow!20 Chain Construction |
-[…]
-orange!20 Compound Conclusion |
-[…]
-mylightblue Concatenation Strategy |
-[…]
-blue!20 Symbolic Strategy |
-Correct Answer |
-Sound Reasoning |
-Zephyr-7B- |
-Mistral-7B-Instruct |
-LLaMA-2-7B |
-LLaMA-2-13B |
+Statements: 1. White xor black. 2. Black xor pink. 3. Pink iff gray. Conclusion: If white then gray.
+Statements: 1. Brown iff orange. 2. Orange xor yellow. 3. Yellow iff green. Conclusion: If brown then green.
+Statements: 1. Green iff purple. 2. If purple then gray. 3. Gray xor yellow. Conclusion: Green xor yellow.
+Statements: 1. Red xor maroon. 2. Maroon xor yellow. 3. Yellow iff orange. Conclusion: If maroon then orange.
+Statements: 1. Purple iff yellow. 2. Yellow iff blue. 3. Blue xor orange. Conclusion: Purple xor orange.
+Statements: 1. Gray iff yellow. 2. Yellow xor olive. 3. Olive iff black. Conclusion: If yellow then black.
+Statements: 1. Blue iff red. 2. Red xor white. 3. White iff pink. Conclusion: If not blue then pink.
+Statements: 1. Olive xor brown. 2. Brown iff gray. 3. Gray xor maroon. Conclusion: If not olive then maroon.
+Statements: 1. Purple iff blue. 2. Blue iff olive. 3. Olive xor red. 4. Red xor green. Conclusion: If purple then green.
+Statements: 1. Brown iff yellow. 2. Yellow xor green. 3. Green iff purple. 4. Purple iff olive. Conclusion: If brown then olive.
+Statements: 1. Red iff maroon. 2. Green xor olive. 3. Maroon iff green. 4. Olive xor brown. Conclusion: If red then brown.
+Statements: 1. Blue iff brown. 2. White iff green. 3. Brown xor white. 4. Green iff purple. Conclusion: If blue then purple.
+| Supposition Following | Chain Construction | Compound Conclusion | Concatenation Strategy | Symbolic Strategy | Correct Answer | Sound Reasoning | |
+| Zephyr-7B- | |||||||
+| Mistral-7B-Instruct | |||||||
+| LLaMA-2-7B | |||||||
+| LLaMA-2-13B | |||||||
 | LLaMA-2-70B |
@@
 We report further details about the models used in this study in Table 3. In particular, we provide information about the number of parameters, context length and fine-tuning procedure for each model.
+| Model | Base Model | Parameters | Context Length | Tokens | Fine-tuning |
+| Zephyr-7B- | Mistral | 7B | 8192 tokens | - | dSFT, AIF |
+| Mistral-7B-Instruct | Mistral | 7B | 8192 tokens | - | SFT |
+| LLaMA-2-7B-Chat | LLaMA-2 | 7B | 4K tokens | 2.0T | SFT, RLHF |
+| LLaMA-2-13B-Chat | LLaMA-2 | 13B | 4K tokens | 2.0T | SFT, RLHF |
+| LLaMA-2-70B-Chat | LLaMA-2 | 70B | 4K tokens | 2.0T | SFT, RLHF |
 Appendix B Additional Quantitative Results
@@
 Each figure is organized with the problem statement at the top, the model’s response on the lower left, and the annotators’ comments to the lower right.
-For an extensive array of model responses and annotations, we invite readers to explore our data repository, which we publish upon acceptance of this paper.
+For an extensive array of model responses and annotations, we invite readers to explore our data repository at: huggingface.co/datasets/mainlp/inferential_strategies.
+| Model | Supposition Following | Chain Construction | Compound Conclusion | Concatenation Strategy | Symbolic Strategy | Correct Answer | Sound Reasoning |
+| Zephyr-7B- | |||||||
+| Mistral-7B-Instruct | |||||||
+| LLaMA-2-7B | |||||||
+| LLaMA-2-13B | |||||||
+| LLaMA-2-70B |
 C.1 Supposition Following
@@
 In Figure 8 the model approaches problem 9 by assuming the presence of an olive marble in the box, yet inferring disjointed intermediate conclusions that do not aid in solving the problem, thus failing to prove the logical validity of the problem.
-Model |
-[…]
-green!20 Supposition Following |
-[…]
-yellow!20 Chain Construction |
-[…]
-orange!20 Compound Conclusion |
-[…]
-mylightblue Concatenation Strategy |
-[…]
-blue!20 Symbolic Strategy |
-Correct Answer |
-Sound Reasoning |
-Zephyr-7B- |
-Mistral-7B-Instruct |
-LLaMA-2-7B |
-LLaMA-2-13B |
-LLaMA-2-70B |
 C.2 Chain Construction
@@
 The model makes false inferences while employing chain construction, and further errs in its validation through logical calculus.
-Model |
-Base Model |
-Parameters |
-Context Length |
-Tokens |
-Fine-tuning |
-Zephyr-7B- |
-Mistral |
-7B |
-8192 tokens |
-dSFT, AIF |
-Mistral-7B-Instruct |
-Mistral |
-7B |
-8192 tokens |
-SFT |
-LLaMA-2-7B-Chat |
-LLaMA-2 |
-7B |
-4K tokens |
-2.0T |
-SFT, RLHF |
-LLaMA-2-13B-Chat |
-LLaMA-2 |
-13B |
-4K tokens |
-2.0T |
-SFT, RLHF |
-LLaMA-2-70B-Chat |
-LLaMA-2 |
-70B |
-4K tokens |
-2.0T |
-SFT, RLHF |
+Statements: 1.
+There is a blue marble in the box if and only if there is a red marble in the box.
+2.
+Either there is a red marble in the box or else there is a white marble in the box, but not both.
+3.
+There is a white marble in the box if and only if there is a pink marble in the box.
+Conclusion: If there is not a blue marble in the box then there is a pink marble in the box.
 Response:
@@
 2. Now, let’s consider the conclusion: "If there is not a blue marble in the box then there is a pink marble in the box"
+Statements: 1.
+There is a blue marble in the box if and only if there is a red marble in the box.
+2.
+Either there is a red marble in the box or else there is a white marble in the box, but not both.
+3.
+There is a white marble in the box if and only if there is a pink marble in the box.
+Conclusion: If there is not a blue marble in the box then there is a pink marble in the box.
 Response:
@@
 1. Conclusion: "If there is not a blue marble in the box then there is a pink marble in the box."
+Statements: 1.
+There is a purple marble in the box if and only if there is a blue marble in the box.
+2.
+There is a blue marble in the box if and only if there is an olive marble in the box.
+3.
+Either there is an olive marble in the box or else there is a red marble in the box, but not both.
+4.
+Either there is a red marble in the box or else there is a green marble in the box, but not both.
+Conclusion: If there is a purple marble in the box then there is a green marble in the box.
 Response:
@@
 However, Statement 4 tells us that there cannot be both a red marble and a green marble in the box either.
+Statements: 1.
+There is a purple marble in the box if and only if there is a blue marble in the box.
+2.
+There is a blue marble in the box if and only if there is an olive marble in the box.
+3.
+Either there is an olive marble in the box or else there is a red marble in the box, but not both.
+4.
+Either there is a red marble in the box or else there is a green marble in the box, but not both.
+Conclusion: If there is a purple marble in the box then there is a green marble in the box.
 Response:
@@
 My reasoning process for this conclusion is as follows:
+Statements: 1.
+There is a blue marble in the box if and only if there is a red marble in the box.
+2.
+Either there is a red marble in the box or else there is a white marble in the box, but not both.
+3.
+There is a white marble in the box if and only if there is a pink marble in the box.
+Conclusion: If there is not a blue marble in the box then there is a pink marble in the box.
 Response:
@@
 Now, let’s consider the conclusion: "If there is not a blue marble in the box then there is a pink marble in the box"
+Statements: 1.
+There is a brown marble in the box if and only if there is a yellow marble in the box.
+2.
+Either there is a yellow marble in the box or else there is a green marble in the box, but not both.
+3.
+There is a green marble in the box if and only if there is a purple marble in the box.
+4.
+There is a purple marble in the box if and only if there is an olive marble in the box.
+Conclusion: If there is a brown marble in the box then there is an olive marble in the box.
 Response:
@@
 Based on the statements we’ve been given, we can logically deduce the conclusion as follows:
+Statements: 1.
+There is a purple marble in the box if and only if there is a yellow marble in the box.
+2.
+There is a yellow marble in the box if and only if there is a blue marble in the box.
+3.
+Either there is a blue marble in the box or else there is an orange marble in the box, but not both.
+Conclusion: Either there is a purple marble in the box or else there is an orange marble in the box, but not both.
 Response:
@@
 Now, let’s consider the conclusion: "Either there is a purple marble in the box or else there is an orange marble in the box, but not both"
+Statements: 1.
+There is a blue marble in the box if and only if there is a brown marble in the box.
+2.
+There is a white marble in the box if and only if there is a green marble in the box.
+3.
+Either there is a brown marble in the box or else there is a white marble in the box, but not both.
+4.
+There is a green marble in the box if and only if there is a purple marble in the box Conclusion: If there is a blue marble in the box then there is a purple marble in the box.
 Response:
@@
 Let’s think about how we can use the given statements to determine whether this conclusion is true or false.
+Statements: 1.
+There is a purple marble in the box if and only if there is a blue marble in the box.
+2.
+There is a blue marble in the box if and only if there is an olive marble in the box.
+3.
+Either there is an olive marble in the box or else there is a red marble in the box, but not both.
+4.
+Either there is a red marble in the box or else there is a green marble in the box, but not both.
+Conclusion: If there is a purple marble in the box then there is a green marble in the box.
 Response:
@@
 Let’s examine the relationship between the statements:
+Statements: 1.
+There is a purple marble in the box if and only if there is a blue marble in the box.
+2.
+There is a blue marble in the box if and only if there is an olive marble in the box.
+3.
+Either there is an olive marble in the box or else there is a red marble in the box, but not both.
+4.
+Either there is a red marble in the box or else there is a green marble in the box, but not both.
+Conclusion: If there is a purple marble in the box then there is a green marble in the box.
 Response:
@@
 Conclusion: If there is a purple marble in the box then there is a green marble in the box
+Statements: 1.
+Either there is an olive marble in the box or else there is a brown marble in the box, but not both.
+2.
+There is a brown marble in the box if and only if there is a gray marble in the box.
+3.
+Either there is a gray marble in the box or else there is a maroon marble in the box, but not both.
+Conclusion: If there is not an olive marble in the box then there is a maroon marble in the box.
 Response:
@@
 This is an if and only if statement, meaning both conditions are equivalent.
+Statements: 1.
+There is a gray marble in the box if and only if there is a yellow marble in the box.
+2.
+Either there is a yellow marble in the box or else there is an olive marble in the box, but not both.
+3.
+There is an olive marble in the box if and only if there is a black marble in the box.
+Conclusion: If there is a yellow marble in the box then there is a black marble in the box.
 Response:
@@
 5. From the third statement, we know that if there is an olive marble in the box, then there is a black marble in the box.
+Statements: 1.
+There is a green marble in the box if and only if there is a purple marble in the box 2.
+If there is a purple marble in the box then there is a gray marble in the box.
+3.
+Either there is a gray marble in the box or else there is a yellow marble in the box, but not both.
+Conclusion: Either there is a green marble in the box or else there is a yellow marble in the box, but not both.
 Response:
@@
 2. Now, let’s consider the conclusion: "Either there is a green marble in the box or else there is a yellow marble in the box, but not both"
+Statements: 1.
+There is a gray marble in the box if and only if there is a yellow marble in the box.
+2.
+Either there is a yellow marble in the box or else there is an olive marble in the box, but not both.
+3.
+There is an olive marble in the box if and only if there is a black marble in the box.
+Conclusion: If there is a yellow marble in the box then there is a black marble in the box.
 Response: